OpenStudy (anonymous):

THe points (3,4,7) and (7,0,9) are endpoints of a diameter of a sphere. a) find the center of the sphere b) find the radius of the sphere c) write a equation for the sphere

OpenStudy (anonymous):

Good question :p are you in calculus 3? This was day 1 material

OpenStudy (anonymous):

do you do the midpoint to find the center of the sphere?

OpenStudy (anonymous):

if so the midpoint would be (3+7/2, 4+0/2, 7+9/2)=(5, 2, 8)

OpenStudy (anonymous):

then the radius would be the distance from any point to the center

OpenStudy (anonymous):

to find the equation you take a point x,y,z and find the distance to the center (5.2.8)

OpenStudy (anonymous):

that would show up as (x-5)^2+(y-2)^2+(z-8)^2=r^2

OpenStudy (anonymous):

and if you need any more help feel free to pm

OpenStudy (anonymous):

are the answers: a) (5,2,8) b) sqrt(29) c) Equation ==> (x-5)^2 + (y-2)^2 + (z-8)^2 = sqrt(29)

OpenStudy (anonymous):

@jefftheloveableguy

OpenStudy (anonymous):

no I got the radius to be 3

OpenStudy (anonymous):

oh..really? hmnmn...weird...so the radius is 3?

OpenStudy (anonymous):

to find the distance between two points in space you use sqrt[(x-x1)^2+(y-y1)^2+(z-z1)^2)}

OpenStudy (anonymous):

so...the radius is 3.

OpenStudy (anonymous):

yep